OpenStudy (anonymous):

i need help withthis multi-step equation.-2(x-9)=-24

OpenStudy (anonymous):

a) divide both sides by \(-2\) to get \[x-9=12\] then aad 9

OpenStudy (anonymous):

make that "add 9"

OpenStudy (anonymous):

x-9= (-24) / (-2) x= 24/2 + 9 x= 12+9 x=21